ELECTRONIC SPECTROSCOPY OF JET-COOLED HCP^{+}

\underline{FUMIE X. SUNAHORI}, XIAOPENG ZHANG, AND DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Laser-induced fluorescence spectra of jet-cooled HCP^{+} and DCP^{+} have been obtained with the pulsed discharge technique using HCP/DCP and argon precursor mixtures. Most of the excited state vibrational fundamentals have been observed and a set of vibrational constants obtained. High resolution spectra of the ^{2}Π_{3/2} components of the 0_{0}^{0} bands of both isotopomers have been recorded, and these spectra show resolved phosphorus hyperfine structure which allowed the determination of the excited state Fermi contact parameter. The B values were used to obtain the ground and excited state effective geometric parameters as r′′_{0}(CH) = 1.077(2) Å, r′′_{0}(CP) = 1.6013(3) Å, r′_{0}(CH) = 1.082(2) Åand r′_{0}(CP) = 1.5331(3) Å. The ground state vibrational energy levels reported in the literature were fitted to a Renner-Teller hamiltonian that included the effects of angular momentum coupling as well as spin-orbit, vibrational anharmonicity and Fermi resonance effects.